What would the format be for a problem like this :a collection of nickels and quarters amounts to $2.60. There are 16 coins in all. How many of each coin are there?
Answer by stanbon(75887)

What would the format be for a problem like this :a collection of nickels and quarters amounts to $2.60. There are 16 coins in all.
How many of each coin are there?
----------------
Equations:
n + q = 16 coins
5n + 25q = 260 cents
---------------------------
Divide thru the lower equation by 5:
n + q = 16
n + 5q = 52
-----
Subtract and solve for "q":
4q = 36
q = 9 (# of quarters)
-----
Solve for "n":
n + q = 16
n + 9 = 16
n = 7 (# of nickels)
===========================
cheers,
Stan H.
==============